OpenStudy (anonymous):

5x + 2y = -15 2x - y = -6 linear equation sub. method

OpenStudy (anonymous):

x=-3 y=0

OpenStudy (whpalmer4):

5x + 2y = -15 2x - y = -6 the second equation can easily be rearranged to give y = 2x + 6, then we substitute that in the first equation: 5x + 2(2x + 6) = -15 5x + 4x + 12 = -15 9x = -27 x = -3 then y = 2(-3)+6 = 0 from the second equation.